OpenStudy (anonymous):

A ball is thrown upward from the ground with an initial speed of 25 m/s, at the same instant, a ball is dropped from rest from a building 15 m high. After how long will the balls be at the same height?

OpenStudy (anonymous):

use equation \[y -y_{o}=\frac{ 1 }{ 2 }g(t -t o)^2 +v _{oy}(t-t o)\] t1 = t2

OpenStudy (anonymous):